# Evaluation and Facilitation of Online Discussions in the LLM Era: A Survey

### v2

@@ 7 Facilitation Datasets
In this section, we provide an overview of the most prominent datasets for online facilitation, considering their sizes and their relevance to core facilitation tasks. We propose the following new taxonomy of facilitation datasets: Conversation Derailment datasets, where the task is to predict when a conversation escalates, therefore requiring facilitator intervention; and Facilitator Interventions datasets, which include comments by facilitators in active discussions, sometimes annotated with the tactics employed. Some datasets contain information that can be used in multiple tasks. An overview of the surveyed datasets and their categories in our taxonomy can be found in Table 1.
Name Task Size Content Wikipedia Disputes De Kock and Vlachos (2021) Conversation Derailment D Includes annotations for several ‘dispute tactics.’ WikiConv Hua et al. (2018) Facilitator Interventions D Includes moderation meta-data such as comment edits and deletions. Conversations Gone Awry Zhang et al. (2018) Conversation Derailment D Predicts derailment by analyzing rhetorical tactics, human-annotated. Chang and Danescu-Niculescu-Mizil (2019) (1) Conversation Derailment D Extends the ‘Conversations Gone Awry’ dataset. Chang and Danescu-Niculescu-Mizil (2019) (2) Conversation Derailment D Based on the r/ChangeMyView subreddit. Park et al. (2012) Conversation Derailment Facilitator Interventions C Comprised of 4 datasets. Includes 19 intervention types belonging to 7 moderator roles. RegulationRoom Falk et al. (2021) Conversation Derailment C Extends the dataset of Park et al. (2012). DeliData Karadzhov et al. (2021) Facilitator Interventions D Group discussions, includes task-oriented quality measure which may be used to approximate discussion quality. Wiki-Tactics De Kock et al. (2022) Facilitator Interventions D Based on Wikipedia Disputes, includes moderation action metadata such as comment edits and deletions. UMOD Falk et al. (2024) Facilitator Interventions C Based on the r/ChangeMyView subreddit, annotated for facilitation tactics and AQ. Fora Schroeder et al. (2024) Facilitator Interventions D Original dataset revolving around experience-sharing, annotated for facilitation tactics.
@@ 11 Ethical Considerations
Although AI and LLMs in particular can be effectively used as discussion facilitators, offering dynamic, responsive discussion support, their deployment must meet strict transparency, safety, and accountability standards, especially for high-risk applications, as stated in the EU AI Act.111https://digital-strategy.ec.europa.eu/en/policies/regulatory-framework-ai For example, a person or minority group may have been unfairly disadvantaged in an AI-enhanced deliberation. It is also necessary for the users to be aware that they are interacting with AI facilitators. Ideally the consent of the users should be sought before using any sort of AI-enhanced discussion platform.
Even if LLMs facilitators eventually achieve a high level of autonomy, it is advisable to maintain human oversight. Keeping a human-in-the-loop ensures greater transparency and enables effective error prevention, detection, and correction.
@@ Appendix C Terminology Background
Here we explain how the surveyed articles were selected. We also explain our reasoning for choosing and disambiguating certain terms (see §2). The definitions of the terms can be found in Table 3.
@@
“Moderation”, as a term, is more common in Computer Science and NLP, while facilitation is prevalent in Social Sciences Vecchi et al. (2021); Kaner et al. (2007); Trenel (2009). Moderators enforce rules and ensure orderly interactions, usually with the threat of disciplinary action, though they can also act as community leaders Falk et al. (2024); Seering (2020); eRulemaking Initiative (2017). Facilitators, on the other hand, guide discussions, promote participation, and structure dialogue, particularly in online deliberation and education platforms Asterhan and Schwarz (2010). Despite these distinctions, the terms are sometimes used interchangeably Cho et al. (2024); Park et al. (2012); Kim et al. (2021), while it is also common for moderators to use facilitation tactics eRulemaking Initiative (2017); Park et al. (2012); Kim et al. (2021); Cho et al. (2024); Schluger et al. (2022).
@@
There is little to no consensus on how to properly define terms such as “discussion” and “dialogue” Russmann and Lane (2016); Goñi (2024). In this section, we attempt to disambiguate the use of such terms for the purposes of our survey and based on the existing related work. First, our study focuses on discussions, a broader term encompassing various informal and formal exchanges, including online discussions in fora Russmann and Lane (2016), with which we are mainly concerned. In contrast, dialogue refers to collaborative interactions in which participants work toward a shared understanding and alignment Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024). Studies on dialogue emphasize its cooperative nature, aiming for mutual insight rather than competition Bawden (2021). Dialogue can also refer to dialogue systems, a major NLP sub-area, traditionally including both task-oriented dialogues and casual conversation (Eliza-like)222http://web.njit.edu/~ronkowit/eliza.html “chatbots” Liu et al. (2023); Sun et al. (2021).
@@ Appendix F Online Discussion Example with Color-coded Politeness Markers
This table highlights key politeness-related linguistic features such as hedging, personal references, sentiment, and direct questions. These features are essential in the context of facilitation, where the goal is to guide conversations constructively, maintain safety, and foster mutual understanding. By identifying these elements, the facilitator (human or automatic) can better interpret the tone, intent, and emotional weight of each utterance. For example, detecting hedging or positive sentiment can guide the model to adopt a more collaborative tone, while recognizing negative sentiment or accusatory second-person references may prompt it to de-escalate tension and encourage constructive dialogue.

### v3

@@ 7 Facilitation Datasets
In this section, we provide an overview of the most prominent datasets for online facilitation, considering their sizes and their relevance to core facilitation tasks. These datasets can be used for analyzing the behavior of human facilitators and the reactions of the participants, investigating the existing taxonomies (e.g., ones presented in §5) or as training data for human and LLM facilitators.
Due to the low number of such datasets in literature Chen et al. (2025), the entries presented in this section straddle various domains adjacent to online facilitation. Hence, we propose the following new taxonomy of facilitation datasets: Conversation Derailment datasets, where the task is to predict when a conversation escalates, therefore requiring facilitator intervention; and Facilitator Interventions datasets, which include comments by facilitators in active discussions, sometimes annotated with the tactics employed. Some datasets contain information that can be used in multiple tasks.111Despite its designation, the ‘WikiDisputes’ dataset does include information about facilitators. We consider it solely a ‘Conversation Derailment’ dataset because facilitator interventions only constitute of its comments. An overview of the surveyed datasets and their categories in our taxonomy can be found in Table 1.
Name Task Size Content WikiDisputes De Kock and Vlachos (2021) Conversation Derailment D Includes annotations for several ‘dispute tactics.’ Wiki-Tactics De Kock et al. (2022) Conversation Derailment Facilitator Interventions D Based on Wikipedia Disputes, includes moderation action metadata such as comment edits and deletions. WikiConv Hua et al. (2018) Facilitator Interventions D Includes moderation meta-data such as comment edits and deletions. Conversations Gone Awry Zhang et al. (2018) Conversation Derailment D Predicts derailment by analyzing rhetorical tactics, human-annotated. Chang and Danescu-Niculescu-Mizil (2019) (1) Conversation Derailment D Extends the ‘Conversations Gone Awry’ dataset. Chang and Danescu-Niculescu-Mizil (2019) (2) Conversation Derailment D Based on the r/ChangeMyView subreddit. Park et al. (2012) Conversation Derailment Facilitator Interventions C Comprised of 4 datasets. Includes 19 intervention types belonging to 7 moderator roles. RegulationRoom Falk et al. (2021) Conversation Derailment C Extends the dataset of Park et al. (2012). UMOD Falk et al. (2024) Facilitator Interventions C Based on the r/ChangeMyView subreddit, annotated for facilitation tactics and AQ. Fora Schroeder et al. (2024) Facilitator Interventions D Original dataset revolving around experience-sharing, annotated for facilitation tactics. WHoW Chen et al. (2025) Facilitator Interventions C Dataset derived from TV debates and radio panels, annotated for facilitation tactics. L2Moderator Gao et al. (2025) Facilitator Interventions 17 D (16.5 hours of transcripts) Facilitated online discussions for ESL speakers.
@@ 11 Ethical Considerations
Although AI, and LLMs in particular, can be effectively used as discussion facilitators, offering dynamic, responsive discussion support, their deployment must meet strict transparency, safety, and accountability standards, especially for high-risk applications, as stated in the EU AI Act.222https://digital-strategy.ec.europa.eu/en/policies/regulatory-framework-ai For example, a person or minority group may have been unfairly disadvantaged in an AI-enhanced deliberation. It is also necessary for the users to be aware that they are interacting with AI facilitators. Ideally, the consent of the users should be sought before using any sort of AI-enhanced discussion platform.
Even if LLMs facilitators eventually achieve a high level of autonomy, it is advisable to maintain human oversight. Keeping a human-in-the-loop approach ensures greater transparency and enables effective error prevention, detection, and correction.
@@ Appendix C Terminology Background
Here, we explain our reasoning for choosing and disambiguating certain terms (see §2). The definitions of the terms can be found in Table 3.
@@
“Moderation”, as a term, is more common in Computer Science and NLP, while facilitation is prevalent in Social Sciences Vecchi et al. (2021); Kaner et al. (2007); Trenel (2009). Moderators enforce rules and ensure orderly interactions, usually with the threat of disciplinary action, though they can also act as community leaders Falk et al. (2024); Seering (2020); eRulemaking Initiative (2017). Facilitators, on the other hand, guide discussions, promote participation and structured dialogue, particularly in online deliberation and education platforms Asterhan and Schwarz (2010). Despite these distinctions, the terms are sometimes used interchangeably Cho et al. (2024); Park et al. (2012); Kim et al. (2021), while it is also common for moderators to use facilitation tactics eRulemaking Initiative (2017); Park et al. (2012); Kim et al. (2021); Cho et al. (2024); Schluger et al. (2022).
@@
There is little to no consensus on how to properly define terms such as “discussion” and “dialogue” Russmann and Lane (2016); Goñi (2024). In this section, we attempt to disambiguate the use of such terms for the purposes of our survey and based on the existing related work. First, our study focuses on discussions, a broader term encompassing various informal and formal exchanges, including online discussions in fora Russmann and Lane (2016), with which we are mainly concerned. In contrast, dialogue refers to collaborative interactions in which participants work toward a shared understanding and alignment Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024). Studies on dialogue emphasize its cooperative nature, aiming for mutual insight rather than competition Bawden (2021). Dialogue can also refer to dialogue systems, a major NLP sub-area, traditionally including both task-oriented dialogues and casual conversation (Eliza-like)333http://web.njit.edu/~ronkowit/eliza.html “chatbots” Liu et al. (2023); Sun et al. (2021).
@@ Appendix F Online Discussion Example with Color-coded Politeness Markers
Table 5 highlights key politeness-related linguistic features such as hedging, personal references, sentiment, and direct questions. These features are essential in the context of facilitation, where the goal is to guide conversations constructively, maintain safety, and foster mutual understanding. By identifying these elements, the facilitator (human or automatic) can better interpret the tone, intent, and emotional weight of each utterance. For example, detecting hedging or positive sentiment can guide the model to adopt a more collaborative tone, while recognizing negative sentiment or accusatory second-person references may prompt it to de-escalate tension and encourage constructive dialogue.